OpenStudy (anonymous):

2x + 3y = 2 x + 6y = 4 I need help.

OpenStudy (anonymous):

Can you use matrices?

OpenStudy (anonymous):

2x + 3y = 2 -(1) x + 6y = 4 - (2) (1)x2 -(2) 2(2x + 3y ) - ( x + 6y) = 2x2 -4 4x -x =0 3x=0 x=0 y = [4 -(0)] /6 = 4/6 =2/3